# ActAlign: Zero-Shot Fine-Grained Video Classification via Language-Guided Sequence Alignment

### v1

@@ 3.5 Dynamic Time Warping
Real-world footage often contains abrupt scene changes or irrelevant frames (e.g., replays, advertisements) that introduce noise into $Z_{i}$. To mitigate this, we apply a 1D moving-average filter of width $w$ across the temporal dimension, as defined in Eq. 7:
@@
Let the smoothed visual embeddings for video $V_{i}$ be $\widetilde{Z}_{i}=[\widetilde{\mathbf{z}}_{i}^{1},\dots,\widetilde{\mathbf{z}}_{i}^{T_{i}}]$ and the sub-action sequence embeddings for class $c_{m}$ be $U_{m}=[\mathbf{u}_{m,1},\dots,\mathbf{u}_{m,K_{m}}]$. We first compute the raw cosine similarity matrix as shown in Eq. 8:
@@
Given the Affinity matrix $\hat{A}^{(m,i)}$ (defined in Eq. 9), we seek a warping path $W^{(m,i)}=\{(k_{1},t_{1}),\dots,(k_{L},t_{L})\}$ that maximizes cumulative similarity under monotonicity and continuity constraints, as formulated in Eq. 10:
@@
s.t. W is a valid warping path between [$\textit{W is a valid warping path between }[1,K_{m}]\textit{ and }[1,T_{i}].$ (10)
We solve this using dynamic programming, as defined in Eq. 11:
@@
with the base case $D_{0,*}=D_{*,0}=-\infty$. The final alignment score is $\max_{k,t}D_{k,t}$, and backtracking recovers the optimal warping path $W^{(m,i)}$.
@@
Upon obtaining the optimal warping path $W^{(m,i)}$ between the smoothed visual sequence $\widetilde{Z}_{i}$ and the sub-action sequence $U_{m}$ for candidate class $c_{m}$, we compute the raw alignment score as defined in Eq. 12:
@@
where $\hat{A}^{(i,m)}$ is the affinity matrix introduced in Eq. 9. To mitigate the bias toward longer warping paths (which can accumulate higher raw scores), we normalize $\gamma_{i,m}$ by the path length, resulting in the average alignment score (Eq. 13):
@@
where $|W^{(i,m)}|$ is the number of matched frame–subaction pairs. Since the similarity values in $\hat{A}^{(i,m)}$ lie in $[0,1]$ (due to the sigmoid in Eq. 9), the normalized alignment score $\hat{\gamma}_{i,m}$ also lies in the range $[0,1]$.
Finally, we predict the class whose sub-action sequence best aligns—on average—with the observed video frames. This is done by selecting the class with the highest normalized alignment score (Eq. 14):
@@ 4 Experiment
Prompt Type Avg. Subactions Avg. Subactions / Domain Avg. Words / Subaction
Short-Fixed ($T=1$) $10.00\pm 0.00$ $10.00\pm 0.00$ $2\pm 0.03$
Context-Rich ($T=1$) $4.94\pm 0.86$ $5.01\pm 0.62$ $13.68\pm 2.78$
Context-Rich ($T=0.2$) $5.04\pm 0.87$ $5.09\pm 0.63$ $13.13\pm 2.55$
Table 1: Linguistic complexity of sub-action scripts generated by different prompting strategies. Context-rich prompts yield fewer but more descriptive sub-actions with significantly higher word counts compared to the Short-Fixed strategy. Lower temperature slightly increases consistency without sacrificing expressiveness.
Method #Param Top-1 (%) Top-2 (%) Top-3 (%)
Random (10 Trials) - $20.81$ $42.04$ $62.50$
Human Evaluation (Oracle) - $61.64$
mPLUG-Owl-Video (Ye et al. 2023) $7B$ $19.49$
VideoChat2 (KunChang Li 2023) $7B$ $21.27$
VideoLLaMA (Zhang, Li, and Bing 2023) $8B$ $22.71$
LLaVA-Next-Video (Zhang et al. 2024) $7B$ $22.90$
Qwen2–VL (Wang et al. 2024b) $7B$ $30.24$
X-CLIP-L/14-16F (Ni et al. 2022) $0.6B$ $16.26$ $33.74$ $49.89$
SigLIP–so400m (Zhai et al. 2023) (mean-pool) $0.9B$ $22.94$ $42.20$ $63.70$
+ DTW Alignment (Ours) $0.9B$ $23.27$ $45.66$ $67.37$
ActAlign (Ours) $0.9B$ $30.51$ $54.34$ $71.05$
Table 2: Zero-shot classification results on ActionAtlas (Salehi et al. 2024) under context-rich (T=0.2) prompting. ActAlign achieves state-of-the-art Top-1, Top-2, and Top-3 accuracy, outperforming all baselines and billion-parameter video–language models without any video–text supervision. These results highlight the effectiveness of structured sub-action alignment over flat representations such as mean-pooling and the open-set recognition capability of image-text models.
@@ 4.1 Experimental Setup
ActionAtlas Dataset
We construct SubActionAtlas upon the ActionAtlas benchmark, which— to the best of our knowledge—remains the most challenging dataset for fine-grained action recognition. It comprises 898 YouTube clips across 56 sports and 558 unique actions. Leveraging its rich action diversity and fine-grained complexity, we adopt ActionAtlas as the foundation for our zero-shot sequence-alignment evaluation. For each clip $V_{i}$, we retain its multiple-choice candidate set $\{c_{i,1},\dots,c_{i,M_{i}}\}$ but replace each high-level label with an LLM-generated sub-action sequence. This extension transforms ActionAtlas into a temporally grounded video–language alignment task, while preserving its original difficulty (human Top-1 accuracy: 61.64%).
@@
We use SigLIP–so400m (patch size 14, $d=384$,  878M parameters). We apply a moving-average smoothing window of $w=30$ frames ($\approx$1s @30 fps) to reduce transient noise and emphasize consistent motion patterns. All experiments run on a single NVIDIA RTX A5000 GPU (25 GB).
Zero‐Shot Setup
Following the zero‐shot protocol, no example videos or sub-action sequence from these classes are used for any training or tuning. For each sample, we decompose each candidate action class $c_{m}$ into its sub-action sequence $S_{m}$ via LLM prompting, then align video frames $V_{i}$ against each sub-action sequence to compute the normalized alignment score $\hat{\gamma}_{i,m}$.
@@ 4.2 Experimental Result
Zero-Shot Baseline Comparisons
We first establish a random prediction baseline, followed by a zero-shot SigLIP baseline using mean-pooled frame embeddings, following ViFi-CLIP (Rasheed et al. 2023). Each video $V_{i}$ is represented by $\bar{\mathbf{z}}_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbf{z}_{i}^{t}$, which is compared via cosine similarity to each class name embedding $\phi_{t}(c_{j})$.
We further compare against open-source video–language models, including entries from the ActionAtlas leaderboard and fine-tuned CLIP variants. Despite using no video–text supervision, our method improves over the SigLIP baseline by 7% Top-1 and 12% Top-2 accuracy, outperforming all baselines and billion-parameter video–language models with $\sim$8$\times$ fewer parameters.
As shown in Table 2, these gains are driven by our sub-action-based temporal alignment, which offers a discriminative representation than frame-level pooling alone.
@@
We ablate each component of ActAlign on SubActionAtlas (Table 3), starting from a mean-pooled SigLIP baseline. Adding DTW alignment introduces temporal structure and yields consistent gains. Context augmentation—injecting domain context (e.g., “Sprint start” $\rightarrow$ “Sprint start in basketball”)—produces the largest boost by reducing semantic ambiguity. Signal smoothing offers a modest but complementary improvement by reducing frame-level noise and clarifying action boundaries.
Configuration Top-1 (%) Top-2 (%) Top-3 (%) SigLIP (Zhai et al. 2023) (mean-pool) $22.94$ $42.20$ $63.70$ + DTW Alignment $25.72$ $45.99$ $66.26$ + Context Augmentation $30.07$ $52.67$ $70.49$ + Signal Smoothing ($w=30$) $30.29$ $53.01$ $70.27$
Table 3: Ablation results under context-rich (T=1.0) prompting (see Table 1). DTW alignment introduces temporal matching, context augmentation reduces sub-action ambiguity, and signal smoothing mitigates frame-level noise.
@@
We evaluate three prompt strategies for generating sub-action scripts using GPT-4o (see Table 1), keeping all other components fixed:
  • Short-Fixed ($T=1.0$): Prompts GPT-4o to generate exactly 10 terse (2–3 word) sub-actions per class using a fixed structure.
  • Context-Rich ($T=1.0$): Produces variable-length, context-rich sub-action scripts incorporating domain-specific cues (e.g., “Wrestler”, “Rider”) at standard temperature ($T=1.0$).
  • Context-Rich ($T=0.2$): Same as CR1.0 but with lower temperature to encourage more consistent and deterministic outputs.
Prompt Description Top-1 (%) Short‐Fixed 2-word, fixed 10 sub-actions $27.06$ Context‐Rich ($T=1.0$) context-rich, variable-length $30.29$ Context‐Rich ($T=0.2$) context-rich, variable-length (low $T$) 30.51
@@
Table 4 shows the performance of each strategy. The context-rich prompt with domain-specific context and low temperature ($T=0.2$) achieve the highest accuracy. In contrast, short-fixed prompts—lacking sufficient semantic specificity—perform worst. These results highlight that reducing ambiguity in sub-action descriptions directly improves alignment quality and classification performance.
@@
Table 5 shows LLM-generated subaction sequences for two Figure skating tactics under our context-rich ($T=0.2$) prompt. The scripts highlight ordered, salient steps enabling precise temporal matching.
@@
Video Example: Performing flying camel spin in figure skating
Biellmann spin $\times$     (Incorrect)
Flying camel spin $\checkmark$     (Correct)
Figure 4: Comparison of similarity heatmaps and DTW alignment paths for a correct classification (right) versus an incorrect prediction (middle). The correct class exhibits clearer segmentation and higher alignment quality. The sub-action scripts are provided in Table 5.
Figure 5: t-SNE visualization of sub-action embeddings without (left) and with (right) context augmentation under Context-Rich prompting ($T=0.2$). Colors indicate different sport domains.
Error Analysis
We identify two primary failure modes in our framework:
  • Ambiguous sub-actions: When LLM-generated scripts include vague steps (e.g., “move to position”), the similarity matrix becomes sparse, limiting DTW’s ability to discriminate between classes. As shown in Figure 6, context-rich prompts yield high-similarity regions than short-fixed ones. Appendix Table 2 highlights the qualitative difference in sub-action specificity.
  • Global alignment bias: DTW enforces full-sequence alignment, which may fail when the action begins mid-clip or exhibits temporal shifts. Without a local alignment mechanism, early or trailing sub-actions introduce noise. Figure 4 illustrates a temporal shift case that does not harm classification in this instance, but could lead to errors when candidate classes are highly similar.
These findings highlight the need for precise sub-action generation and motivate improvements in action prototype design.
(a) Video Snapshot: Representative frame showing the execution of Varial kickflip underflip body varial in Skateboarding.
(b) Short‐Fixed (incorrect, $\hat{\gamma}=0.35$).
(c) Context‐Rich ($T=0.2$, correct, $\hat{\gamma}=0.81$).
Figure 6: Failure case due to vague sub-actions for the action Varial kickflip underflip body varial. (6(b)) The Short-Fixed prompt yields a noisy similarity map and incorrect prediction. (6(c)) The Context-Rich prompt produces clearer alignment and correct classification. See Appendix Table 2 for scripts.
Sub-action Embedding Structure
Figure 5 shows 2D t-SNE projections of sub-action embeddings with and without context augmentation under context-rich prompts ($T=0.2$). Adding domain-specific cues ({sub-action} in {sport name}) results in tighter, more coherent clusters—indicating better semantic structure and separation.
Effect of Signal Smoothing
@@
(a) Without smoothing
(b) With smoothing ($w=30$)
Figure 7: Similarity matrices before and after applying a moving-average filter ($w=30$). Signal smoothing reduces high-frequency noise and enhances transition between sub-actions.
@@ 5 Conclusion
We show that contrastive image–text models, even when used with simple mean-pooling, establish a surprisingly strong baseline for open-set fine-grained video classification. To fully leverage this capability in extremely fine-grained settings, we propose ActAlign, a novel zero-shot framework that revisits the classic Dynamic Time Warping algorithm to cast video classification as a sequence-matching problem. By aligning video frames with LLM-generated sub-action scripts, ActAlign introduces temporal structure into contrastive models without requiring any video–text training or fine-tuning. Evaluated on the highly challenging ActionAtlas benchmark, our method achieves state-of-the-art performance, outperforming both CLIP-style baselines and billion-parameter video–language models. These results underscore the value of structured language priors and classical alignment in unlocking the open-set recognition potential of vision–language models for fine-grained video classification.
Future Work.
The effectiveness of ActAlign remains limited by the quality of LLM-generated sub-action scripts. Future work includes exploring improved prompting strategies and local alignment techniques. Given its general design, ActAlign can be integrated with any vision–language model to advance zero-shot fine-grained video understanding across other domains.

### v2

@@ 3.5 Dynamic Time Warping
Real-world footage often contains abrupt scene changes or irrelevant frames (e.g., replays, advertisements) that introduce noise into $Z_{i}$. To mitigate this, we apply a 1D moving-average filter of width $w$ across the temporal dimension:
@@
Let the smoothed visual embeddings for video $V_{i}$ be $\widetilde{Z}_{i}=[\widetilde{\mathbf{z}}_{i}^{1},\dots,\widetilde{\mathbf{z}}_{i}^{T_{i}}]$ and the sub-action sequence embeddings for class $c_{m}$ be $U_{m}=[\mathbf{u}_{m,1},\dots,\mathbf{u}_{m,K_{m}}]$. We first compute the raw cosine similarity matrix:
@@
Given the affinity matrix $\hat{A}^{(m,i)}$ for class $c_{m}$ and video $V_{i}$ (defined in Eq. 9), we seek a warping path $W^{(m,i)}=\{(k_{1},t_{1}),\dots,(k_{L},t_{L})\}$ that maximizes cumulative similarity under monotonicity and continuity constraints:
@@
s.t. $\textit{W is a valid warping path between }[1,K_{m}]\textit{ and }[1,T_{i}].$ (10)
We solve this using dynamic programming by selecting the path with highest alignment (i.e. similarity value) at each step:
@@
with the base case $D_{0,*}=D_{*,0}=-\infty$. The final alignment score is $\max_{k,t}D_{k,t}$, and backtracking recovers the optimal warping path $W^{(m,i)}$ for sub-action sequence of class $c_{m}$ and video $V_{i}$.
@@
We compute the raw alignment score, as defined in Eq. 12, as the sum of the similarity values of the optimal warping path $W^{(m,i)}$ obtained in Eq. 10.
@@
$\hat{A}^{(i,m)}$ is the affinity matrix introduced in Eq. 9. To mitigate the bias toward longer warping paths (which can accumulate higher raw scores), we normalize $\gamma_{i,m}$ by the path length, resulting in the average alignment score:
@@
where $|W^{(i,m)}|$ is the number of matched frame–subaction pairs. Since the similarity values in $\hat{A}^{(i,m)}$ lie in [0,1][0,1][ 0 , 1 ] (due to the sigmoid in Eq. 9), the normalized alignment score $\hat{\gamma}_{i,m}$ also lies in the range [0,1][0,1][ 0 , 1 ].
Finally, we predict the class whose sub-action sequence best aligns—on average—with the observed video frames. This is done by selecting the class with the highest normalized alignment score:
@@
Method #Param Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$
Random (10 Trials) - 20.8120.8120.81 42.0442.0442.04 62.5062.5062.50
Human Evaluation (Oracle) - 61.64 - -
mPLUG-Owl-Video (Ye et al. 2023) $7B$ 19.4919.4919.49 - -
VideoChat2 (Li et al. 2023) $7B$ 21.2721.2721.27 - -
VideoLLaMA (Zhang, Li, and Bing 2023) $8B$ 22.7122.7122.71 - -
LLaVA-Next-Video (Zhang et al. 2024) $7B$ 22.9022.9022.90 - -
Qwen2–VL (Wang et al. 2024b) $7B$ 30.2430.2430.24 - -
X-CLIP-L/14-16F (Ni et al. 2022) $0.6B$ 16.2616.2616.26 33.7433.7433.74 49.8949.8949.89
SigLIP–so400m (Zhai et al. 2023) (mean-pool) $0.9B$ 22.9422.9422.94 42.2042.2042.20 63.7063.7063.70
+ DTW Alignment (Ours) $0.9B$ 25.7225.7225.72 45.9945.9945.99 66.2666.2666.26
ActAlign (Ours) $0.9B$ $30.29$ $53.01$ $70.27$
Table 1: Zero-shot classification results on ActionAtlas. Our method achieves state-of-the-art Top-1, Top-2, and Top-3 accuracy, outperforming all baselines and billion-parameter video–language models without any video–text supervision. These results highlight the open-set recognition capability of image-text models and the effectiveness of structured sub-action alignment over flat representations such as mean-pooling. We use context-rich prompting strategy.
@@ 4.1 Experimental Setup
Dataset
Our approach is dataset- and domain-agnostic, as it relies solely on language-driven sub-action generation from LLMs and cross-modal alignment without task-specific supervision. To rigorously evaluate its generality, we adopt ActionAtlas (Salehi et al. 2024)—to our knowledge, the most diverse and challenging benchmark for fine-grained action recognition across various domains. For each video $V_{i}$, we retain its multiple-choice candidate set $\{c_{i,1},\dots,c_{i,M_{i}}\}$, and replace each class label with an LLM-generated sub-action sequence. (Dataset statistics is provided in Appendix.)
@@
We use SigLIP–so400m (patch size 14, $d=384$,  878M parameters). We apply a moving-average smoothing with a fixed window size of $w=30$ frames (1s @30 fps) to reduce transient noise and emphasize consistent motion patterns. All experiments run on a single NVIDIA RTX A5000 GPU (25 GB). Inference consists of embedding extraction followed by DTW alignment over $N$ class candidates, each with $M$ sub-actions, and $T$ video frames—yielding $\mathcal{O}(NMT)$ time. Following the zero‐shot protocol, no example videos or sub-action sequence from these classes are used for any training or tuning. We only use candidate action class names $c_{m}$ to prompt for sub-action generation.
@@ 4.2 Experimental Result
Zero-Shot Comparisons
We begin with a random choice baseline and a SigLIP zero-shot baseline using mean-pooled frame embeddings. Following ViFi-CLIP (Rasheed et al. 2023), each video $V_{i}$ is encoded as $\bar{\mathbf{z}}_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbf{z}_{i}^{t}$, where mean-pooled embeddings $\bar{\mathbf{z}}$ is compared via cosine similarity to each name embedding $\phi_{t}(c_{j})$.
We further compare against open-source video–language models, including entries from the ActionAtlas leaderboard and fine-tuned CLIP variants. Despite using no video–text supervision, our method outperforms the SigLIP baseline by +7% Top-1 and +11% Top-2 accuracy, surpassing all baselines and billion-parameter models with $\sim$8$\times$ fewer parameters. These larger models are often optimized for interactive tasks through extensive instruction tuning, rather than for classifying fine-grained actions. In contrast, image–language models are trained on large-scale image–text pairs, making them particularly well-suited for open-set recognition when paired with temporal alignment mechanisms.
As shown in Table 1, these gains stem from our subaction-level alignment, which enables more discriminative and interpretable classification than global frame pooling method.
@@
We ablate each component of ActAlign on ActionAtlas (Table 2), starting from a mean-pooled SigLIP baseline. Adding DTW alignment introduces temporal structure and yields consistent gains. Context augmentation—injecting domain context (e.g., “Sprint start” $\rightarrow$ “Sprint start in basketball”)—produces the largest boost by resolving semantic ambiguity. Signal smoothing offers a modest but complementary improvement by reducing frame-level noise and clarifying action boundaries.
Configuration Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$ SigLIP (Zhai et al. 2023) (mean-pool) 22.9422.9422.94 42.2042.2042.20 63.7063.7063.70 + DTW Alignment 25.7225.7225.72 45.9945.9945.99 66.2666.2666.26 + Context Augmentation 30.0730.0730.07 52.6752.6752.67 70.4970.4970.49 + Signal Smoothing 30.2930.2930.29 53.0153.0153.01 70.2770.2770.27
Table 2: Ablation results of various design under context-rich prompting. DTW alignment introduces temporal matching, context augmentation reduces sub-action ambiguity, and signal smoothing mitigates frame-level noise.
@@
We evaluate two prompt strategies for generating sub-action scripts using GPT-4o, keeping all other components fixed:
  • Short-Fixed: Prompts GPT-4o to generate exactly 10 terse (2–3 word) sub-actions per class using a fixed structure.
  • Context-Rich: Produces variable-length, context-rich sub-action scripts incorporating domain-specific cues (e.g., “Wrestler”, “Rider”).
Prompt Description Top-1 (%) $\uparrow$ Short‐fixed 2-word, fixed 10 sub-actions 27.0627.0627.06 Context‐rich context-rich, variable-length 30.29
@@
Table 3 shows the performance of each strategy. The context-rich prompt with domain-specific context achieve the highest accuracy. In contrast, short-fixed prompts—lacking sufficient semantic specificity—perform worst. These results highlight that reducing ambiguity in sub-action descriptions directly improves alignment quality and classification performance.
Sub-action Embedding Structure
Figure 3 shows 2D t-SNE projections of sub-action embeddings with and without context augmentation under context-rich prompts. Adding domain-specific cues (<sub-action> in <sport name>) results in tighter, more coherent clusters—indicating better semantic structure and separation.
w/o Context Augmentation
w/ Context Augmentation
Figure 3: t-SNE visualization of sub-action embeddings. Each color corresponds to a sport domain. Augmenting sub-actions with context yields more discriminative clusters and improves textual grounding.
@@
Table 4 shows LLM-generated subaction sequences for two Figure skating tactics under our context-rich prompting. The scripts highlight ordered, salient steps enabling precise temporal matching.
@@
(a) Video: Performing flying camel spin in figure skating.
(b) Biellmann spin $\times$
(c) Flying camel spin $\checkmark$
Figure 4: DTW alignment paths for an incorrect prediction (left) versus a correct classification (right). The correct class exhibits clearer segmentation and higher alignment quality. The sub-action scripts are provided in Table 4.
@@
Signal Smoothing.
@@
(a) w/o smoothing
(b) w/ smoothing
Figure 5: Signal smoothing reduces high-frequency noise and enhances transition between sub-actions. Similarity matrices before and after applying a moving-average filter ($w=30$).
(a) Video: Varial kickflip underflip body varial in Skateboarding.
(b) Short‐fixed ($\times$, $\hat{\gamma}=0.35$).
(c) Context‐rich ($\checkmark$, $\hat{\gamma}=0.81$).
Figure 6: Failure case due to vague sub-actions for the action Varial kickflip underflip body varial. The normalized alignment score $\hat{\gamma}$ reflects overall sequence similarity. (6(b)) The Short-fixed prompt yields a noisy similarity map and incorrect prediction. (6(c)) The Context-rich prompt produces clearer alignment and correct classification. See Appendix for the detailed scripts.
Failure Case Analysis
We observe two primary failure modes in our framework:
  • Ambiguous sub-actions: Vague LLM-generated steps (e.g., “move to position”) result in sparse similarity matrices, weakening DTW’s ability to discriminate between candidates. As shown in Figure 6, context-rich prompts produce clearer alignment regions than short-fixed ones.
  • Global alignment bias: Vanilla DTW enforces full-sequence alignment, which is suboptimal when actions start mid-clip or exhibit temporal shifts. Without a local alignment mechanism, early or trailing sub-actions can introduce noise. Figure 4 shows a temporal shift case that does not affect prediction here but may cause errors in more competitive settings.
These limitations underscore the importance of precise sub-action design and motivate future improvements in alignment robustness and prototype refinement.
@@ 5 Conclusion
We show that contrastive image–language models establish a surprisingly strong baseline for zero-shot fine-grained video classification, even when used with simple mean-pooling. To fully leverage this capability in extremely fine-grained settings, we propose ActAlign, a novel zero-shot, domain-general, and model-agnostic framework that revisits the classic Dynamic Time Warping (DTW) algorithm to cast video classification as a sequence-alignment problem. By aligning video frames with LLM-generated sub-action scripts, ActAlign introduces temporal structure into contrastive models without requiring any video–text training or fine-tuning. Evaluated on the highly challenging and diverse ActionAtlas (Salehi et al. 2024) benchmark, our method achieves state-of-the-art performance, outperforming both CLIP-style baselines and billion-parameter video–language models.
Appendix A Smoothing Analysis
We assess the effect of signal smoothing on model performance. Specifically, we vary the moving-average window size $w\in\{10,20,30,50\}$, applied over the frame-wise SigLip embeddings before alignment.
Effect on Accuracy.
As shown in Figure 7, the Top-1 classification accuracy remains almost stable across a wide range of smoothing widths. This suggests that our method is not overly sensitive to the smoothing parameter.
Figure 7: Top-1 accuracy for different smoothing window sizes $w$. Accuracy is stable across settings, with $w=30$ selected as default.
Choice of Default.
We fix $w=30$, which corresponds to a 1-second temporal window under 30 FPS videos. This setting balances local context preservation with jitter reduction and yields smooth DTW alignments that more cleanly segment sub-actions.
Appendix B Implementation Details and Inference Pipeline
System Configuration
We conduct all experiments on a server running Ubuntu 20.04 with a single NVIDIA RTX A5000 GPU (24GB VRAM) and 256GB RAM. The pipeline is implemented in PyTorch 2.7 with CUDA 12.8. For subaction generation, we use the GPT-4o API.
Model Settings
We use the publicly available SigLIP model from Google (siglip-so400m-patch14-384), which supports variable input resolutions. Images are embedded using a ViT-based vision encoder with a patch size of $\times 14$ 14 and an output embedding dimension of $d=384$. Sub-actions are encoded using the corresponding frozen SigLIP text encoder. All embeddings are $\ell_{2}$-normalized before computing cosine similarities.
Runtime Performance
End-to-end inference over all 898 samples—including affinity matrix computation, signal smoothing, temporal alignment, and classification—completes in 41.95 seconds, with an average processing time of 0.04 seconds per video. Dynamic time warping (DTW) alignment accounts for over 90% of the total runtime.
Inference Pipeline Overview.
Our pipeline proceeds in six modular stages: (1) videos from ActionAtlas are downloaded and clipped using annotated timestamps, with frames stored in a structured NumPy archive; (2) each frame is encoded using SigLip’s vision encoder to obtain per-frame embeddings; (3) GPT-4o generates context-rich sub-action sequences for each candidate label using fixed prompt templates; (4) sub-actions are encoded via SigLip’s text encoder; (5) cosine similarity matrices are computed between visual and semantic sequences, and DTW is applied to compute alignment scores for each candidate; and (6) as a baseline, we compare mean-pooled frame embeddings with text embeddings of class labels using cosine similarity, bypassing alignment. Each step is modular and implemented for fast inference, allowing to easily switch between models and datasets.
Appendix C More Related Works
Image–Language Transfer for Video Tasks. Image–language models pretrained on large-scale image–text pairs exhibit strong open-set recognition capabilities. This motivates their adaptation for video tasks. Common strategies include adding temporal modules (Wang, Xing, and Liu 2022; Ni et al. 2022; Rasheed et al. 2023; Liu, Li, and Yu 2024) or introducing learnable prompts (Ju et al. 2022; Lin et al. 2022). However, as noted by (Rasheed et al. 2023), fine-tuning often compromises the model’s open-set generalization. To preserve zero-shot capabilities, recent work explores training-free adaptations using temporal prompting methods (Ahmad, Chanda, and Rawat 2024; Phan et al. 2024; Yu et al. 2025). For example, TEAR (Bosetti et al. 2025) compares mean-pooled frame embeddings with LLM-generated action descriptions for each class. While effective, these approaches either degrade generalization through fine-tuning or ignore temporal structure via mean-pooling - limiting performance on fine-grained and zero-shot video classification.
Appendix D Prompt Templates and Generation Settings
We use GPT-4o to automatically convert each high-level class name into a sequence of sub-actions. Two prompt styles were explored:
D.1 Short-Fixed Prompt.
This setting produces terse outputs (e.g., “tie arms,” “deliver knee”) that often lack domain-specific context. The descriptions are directly adopted from the orginal ActionAtlas (Salehi et al. 2024) dataset.
Template
Generate a JSON object from the following context.
For each action option given below, create a key in the JSON object with that action name.
The value for each key must be a list of exactly 10 subactions.
Each subaction must be a minimalist, two‐word description.
The descriptions should capture the essential mechanics of the action.
Use the following action options and their descriptions for context:
- <action 1>: <action 1 description>
- <action 2>: <action 2 description>
- <action 3>: <action 3 description>
Output the JSON object accordingly.
D.2 Context-Rich Prompt.
This setting encourages richer descriptions with more related keywords grounded in the activity domain. Context-rich prompting strategy significantly improves textual grounding during DTW alignmnet.
Template
You will output a JSON object. Each key is an action name from the list below, and its value is a list of concise, visually distinctive subaction descriptions performed sequentially.
Do NOT fix the list length—choose however many substeps best break down that trick into discriminative subactions.
Each subaction should be:
 – Self-contained and visually descriptive with rich <domain> keywords and context.
 – Explicitly reference the scene, objects, environment, and motion in the given sport context.
 – Discriminative and concise.
Here are the actions to decompose:
<action 1> in <domain>
<action 2> in <domain>
<action 3> in <domain>
Output ONLY the JSON object, without any additional explanation.
D.3 Examples
Table 5 compares the two prompting strategies for an skateboarding trick example. The corresponding affinity matrices are shown in Figure 6.
Note.
Our experiments demonstrate that context-rich prompts significantly outperform short-fixed ones, as the added context helps disambiguate tactics across sports and yields smoother DTW alignment with video frames.
Prompt Variant Generated Subaction Script
Short‐fixed 1. flip board
2. spin board
3. rotate body
4. catch board
5. kick upward
6. flip reverse
7. control motion
8. hover steady
9. anticipate landing
10. land cleanly
Context‐rich 1. Rider rolls forward on flat ground, knees compressed and eyes focused ahead.
2. Back foot scoops the tail into a backside shuvit while front foot flicks a kickflip off the nose-side edge.
3. As the board begins its varial kickflip rotation, the front foot quickly retracts and then kicks downward under the board’s underside.
4. Board completes a varial kickflip and then reverses flip direction mid-air due to the underflip, creating a complex double-flip motion.
5. Simultaneously, rider initiates a 180-degree body varial, rotating in the same direction as the board’s shuvit.
6. Rider tracks the board’s griptape, completes the body spin, and catches the board with both feet aligned over the bolts.
7. Rider lands smoothly, now facing the opposite direction, and rolls away with balance.
Table 5: Sub-action sequence comparison for the varial kickflip underflip body varial trick in skateboarding using different prompt variants. Figure 6 represents the heatmap comparisons.
Appendix E Dataset Statistics
Dataset Overview.
Table 6 summarizes key statistics of the ActionAtlas (Salehi et al. 2024) dataset, which we extended to enable alignmnet-based classification. The dataset contains 898 video clips spanning 558 unique tactics across 56 sports domains. Each video is paired with 4 to 6 fine-grained candidate actions (5.25 in average per video). Videos vary significantly in length, with a mean of 173 frames (30 FPS), ranging from extremely short clips (3 frames) to multi-second sequences (up to 900 frames). Notably, ActionAtlas is the most diverse and challenging fine-grained action recognition benchmark to the best of our knowledge. It poses a major challenge in fine-grained recognition: human top-1 accuracy is just 61.64%.
Category Statistic Value
Dataset scale Videos 898
Unique fine-grained actions 558
Sport domains 56
Avg. candidate actions per video 5.26 ± 1.27
Avg. candidate actions per domain 33.6
Min–Max candidate actions per domain 4–201
Temporal properties Avg. video length (frames) 173.2 ± 121.3
Min–Max video frames 3–900
Recognition accuracy Human top-1 accuracy (%) 61.64
Table 6: Summary of ActionAtlas dataset statistics.
ActionAtlas Extension.
To enable subaction-level alignment, we augment the original ActionAtlas dataset with sub-action scripts generated by a large language model for each candidate action (see Figure 8). This transforms the task from multiple-choice classification into alignment-based classification without modifying the original videos or labels. The extended dataset supports zero-shot video classification via sub-action sequence matching.
Figure 8: Our pipeline uses an LLM to generate structured sub-action sequences for each fine-grained candidate action in ActionAtlas (Salehi et al. 2024) data sample. This structured representation enables sequence alignment with video frames for zero-shot recognition.
Sub-action Script Statistics.
We generate sub-action sequences using different prompting strategies. As summarized in Table 7, context-rich prompts yield fewer but more descriptive sub-actions, while short-fixed prompts produce uniform-length, terse sequences. These linguistic differences influence the textural grounding, and consequently, sequence alignment.
Prompt Type Avg. Subactions Avg. Subactions per Domain Avg. Words per Subaction
Short-fixed 10.00 ± 0.00 10.00 ± 0.00 2.00 ± 0.03
Context-rich 4.94 ± 0.86 5.01 ± 0.62 13.68 ± 2.78
Table 7: Linguistic complexity of subaction scripts generated by different prompting strategies. Context-rich prompts yield fewer but more descriptive subactions with significantly higher word counts compared to the Short-fixed strategy.